So there is this problem that I have for homework that says "Solve 6(x-6)+6=8(x+2)-8. I'm not sure if I should solve for x. For a problem like this, is using FOIL the best method?
Answer by ankor@dixie-net.com(12706)

"Solve 6(x-6)+6=8(x+2)-8.
:
Just multiply what's inside the brackets, combine like terms
6x - 36 + 6 = 8x + 16 - 8
6x - 30 = 8x + 8
:
Combine the unknowns (x) on the left, and numerical values on the right
6x - 8x = 8 + 30
-2x = 38
:
x has to be positive, multiply equation by -1:
2x = -38
x = -38/2
x = -19
:
:
Check solution, substitute -19 for x in the original equation:
6(x-6) + 6 = 8(x+2)- 8.
6(-19-6) + 6 = 8(-19+2) - 8
6(-25) + 6 = 8(-17) - 8
-150 + 6 = -136 - 8
-144 = -144; equality reigns confirming our solution
:
FOIL is not used in this type of problem